# A flexible and general semi-supervised approach to multiple hypothesis testing

### v1

@@ 5.2 Application to peptide detection
In this section we compare RESET to Adaptive Knockoffs in the peptide detection context using mass spectrometry data and show that (a) the Adaptive Knockoff methods are often computationally infeasible, (b) feasibility aside, RESET is more powerful than Adaptive Knockoff’s default filter, AdaKO GAM, when controlling the FDR and (c) RESET is able to typically discover more peptides than the vanilla FDP-SD when controlling the FDP. We provide justification of Assumption 3 which is required for RESET’s theoretical guarantees in Section S.15.
@@ 5.2.1 HEK293 data
We show that the application of Adaptive Knockoffs is often too slow to be used effectively in high throughput scenarios such as the peptide detection problem. We demonstrate this, using HEK293 data [9], a popular dataset used in benchmarking the development of new software [30, 49, 32] that we downloaded from the Proteomics Idenitification Database (PRIDE) [40]. After data preparation and searching the combined 24 HEK293 spectrum files, we determined the scores $Z_{i}$ for each target peptide and $\tilde{Z}_{i}$ for each decoy peptide using the XCorr score function (see Supplementary Section S.12 for details). Next, we competed each pair by recording the score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and label $L_{i}=1$ if $Z_{i}>\tilde{Z}_{i}$ and $L_{i}=-1$ if $Z_{i}<\tilde{Z}_{i}$ (ties were randomly broken). In addition to the primary score associated with each peptide-spectrum match (PSM) is a set of features, such as the charge state or peptide length. Each winning peptide inherits these features, or side information $\mathbf{x}_{i}$, from its maximally associated PSM, allowing us to define the desired triples $(L,W,\mathbf{x})$. A complete description of the side information used is given in Table S4 of the Supplement.
@@ 5.2.2 PRIDE data
Due to the aforementioned runtime problem, we had to use small datasets when evaluating the power of Adaptive Knockoffs in the peptide detection context. Specifically, we focused on thirteen spectrum files (each with $<$ 33K PSMs) from the PRIDE-20 dataset that we recently used in [18]. In the following results, we only considered searching the spectrum files once against a single target-decoy database since unlike RESET we can only feasibly apply a single application of Adaptive Knockoffs. Data preparation, search settings and details on how each triple $(L,W,\mathbf{x})$ is obtained are given in Supplementary Section S.14 along with the list of side information used in Supplementary Table S5.
@@
Figure 5: Comparison of RESET Ensemble and AdaKO GAM. For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed. The figure shows a boxplot of those average RESET Ensemble discoveries divided by the number of discoveries using AdaKO GAM (left boxplot) and TDC (right boxplot) using the 13 PRIDE-20 datasets (there are total of 13 points for each boxplot).
@@
Figure 5 shows the average number of discoveries reported using 10 applications of RESET Ensemble, each applied at 1% FDR to the same dataset, divided by the number of discoveries reported by AdaKO GAM (left boxplot) and TDC (right boxplot). In each of the 10 applications of RESET, we varied the internal seed which randomly splits the decoys into their training and estimating sets. Clearly, RESET Ensemble is a more powerful alternative than AdaKO GAM and TDC. Interestingly, the relative ratios in the number of discoveries between AdaKO GAM and TDC suggests that AdaKO GAM is reporting less discoveries than the generic TDC.
@@
In a second comparision, we looked at the performance of RESET Ensemble versus AdaKO RF because in this setup it was the most powerful implementation of Adaptive Knockoffs. In Figure 6A, the left boxplot displays the average number of discoveries by RESET Ensemble, divided by the number of discoveries by AdaKO RF. In six of the thirteen PRIDE spectrum files, we found RESET Ensemble to be the superior method, while in the remaining seven spectrum files, AdaKO RF produced more discoveries. In one data set, AdaKO RF discovered 1217 peptides while RESET Ensemble only discovered 1104 peptides, about a 10% increase over RESET Ensemble.
@@
We also assessed the computational times of AdaKO RF in Figure 6B, using both the actual and estimated runtimes as outlined in Section S.13. Reassuringly, our estimated times were less than the actual times taken by each method to complete thus validating our estimated times in the previous section. Although we find that, when analyzing these smaller spectrum sets, AdaKO RF is the most powerful method of Adaptive Knockoffs, and is competitive to RESET Ensemble power-wise, it is again clear that AdaKO RF is generally too slow with some runtimes exceeding multiple days even for these small datasets. In fact, on the dataset where AdaKO RF outperforms RESET Ensemble by about 10%, it takes 3.07 days for AdaKO RF to discover its peptides while RESET Ensemble takes about half a minute to complete. Figure 6C shows that in the worst case, RESET takes roughly 1 minute. The issue of runtimes is especially problematic considering the fact that multiple spectrum files are often analyzed jointly, as in our HEK293 analysis in the previous section.
@@ 5.2.3 FDP control
Lastly, we evaluated RESET Ensemble’s FDP control using the HEK293 and PRIDE-20 spectrum files from the previous sections. As in Section 5.2.1, the joint HEK293 spectrum files were searched against 5 distinct target-decoy databases, where each target-decoy database was generated by concatenating the target database to a decoy database consisting of shuffled target peptides. Similarly, each PRIDE-20 spectrum file was searched against 10 distinct target-decoy databases. We used the coinflip version of FDP-SD [39], which provides a uniform increase in the number of discoveries (at the cost of some randomness in the discovery list, see Supplementary Algorithm S2). Our results using RESET Ensemble and FDP-SD are given in Table 3 and Figure 7.
@@
Analyzing the HEK293 results first, RESET Ensemble is clearly more powerful with an approximately 9-10% increase in the average number of discoveries over FDP-SD at all confidence levels. Analyzing the PRIDE-20 datasets, while RESET Ensemble appears to be generally more powerful, in a few datasets, it fails to make any discoveries at high confidence levels. Upon a closer look, the numbers of discoveries made by FDP-SD in these datasets are relatively small, with the largest average number of discoveries among these datasets being 417. When the number of discoveries is expected to be small and the confidence parameter is high, RESET may yield less power, even if the side information is reasonably informative. We dwell on this fact in the Discussion. Regardless, more often than not RESET Ensemble produces more discoveries and if multiple spectrum files are analyzed jointly, as we do with the HEK293 spectrum files, then the risk of having too few discoveries is reduced.
Figure 7: Comparison of RESET Ensemble and FDP-SD. The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD for each of the PRIDE-20 datasets at a 1% FDP threshold and varying confidence parameters. The averages are taken over 10 applications, one for each of the 10 target-decoy databases used for each PRIDE-20 dataset. The right panel is a ‘zoomed’ in version of the left panel.
The results above consider a canonical ‘narrow’ search. Specfically, during the searching phase, each database peptide and spectrum are only allowed to match if the mass of the database peptide and the mass of the sample peptide generating the spectrum, the ‘spectrum mass’, differ by a small amount. Recently, ‘open’ searching, where each spectrum and peptide can match even if their masses differ by hundreds of Daltons, has become a popular method for detecting peptides [30]. Such a search allows for the detection of peptides whose mass have been altered by post-translational modifications (PTMs). We provide further detail in Supplementary Section S.14. In this setting, side information variables such as dm, which records the mass difference between the peptide and the spectrum, could be highly informative given that some differences will correspond to masses of common PTMs. Thus, it is not surprising that RESET, which can take advantage of such side information, does even better than FDP-SD in this context. Specifically, Figure 8 shows a greater improvement over FDP-SD across all confidence levels when analyzing open searching results.
@@ 6.1 Numerical experiments
In this section, we consider the p-value setting, comparing RESET with the FDR controlling procedures we reviewed in Section 3.1 on a set of simulations introduced by Lei and Fithian [33]. As in the competition setting, each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees. Details of each method used is given in Supplementary Section S.6-S.10.
@@ 6.1.1 Simulation 5: p-value based testing with geometric side information
In this section, we look at Simulation 1 of Lei and Fithian [33]. For each hypothesis, they obtain a p-value $p_{i}$ from a one-sided normal test so that $p_{i}=1-\Phi(z_{i})$ where $z_{i}\sim\mathcal{N}(\mu,1)$ with $\mu=2$ for a false null hypothesis and $\mu=0$ for a true null hypothesis. A total of $m=2500$ hypotheses are used. Similar to Simulation 3 of Section 5.1.3, each hypothesis’ side information is defined as unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from $50\times 50$ equally spaced values in the square $\left[-100,100\right]\times\left[-100,100\right]$. The false null hypotheses correspond to certain regions in the square. They considered three scenarios: (a) when the false nulls form a circle in the middle of the square, (b) when the false nulls form a circle in the top right of the square and (c) when the false nulls form an ellipse. Each of these scenarios is depicted in Figure 9. We used 100 independent runs where in each run, we randomly draw the 2500 p-values as described.
@@
Figure 9: Location of false null hypotheses according to the pair of values $(x,y)$ in the square $\left[-100,100\right]\times\left[-100,100\right]$. The same picture is given in [33].
@@
Figure 10: RESET vs. other methods using two-dimensional side information with p-values. The first row plots the power and the second row plots the estimated FDR for each FDR threshold ranging from 1% to 20%. Each color and shape combination corresponds to a unique method. The three columns correspond to the hypotheses’ relationship to the side-information according to Figure 9. The horizontal segments in the second row of panels mark the corresponding FDR thresholds on the y-axis. For readability, each point is jittered in the horizontal direction.
@@
Figure 10 shows the results of RESET compared with several p-value based methods that we reviewed in Section 3.1. We find that in the first two simulations, (a) and (b), RESET Ensemble is roughly equal best in terms of power at the 5%, 10% and 20% FDR levels, along with AdaPT/AdaPTg GAM. However, in (c), clearly AdaPT/AdaPTg GAM are more powerful than the RESET methods, with RESET Ensemble coming arguably as third in the power ranking at the 5%, 10%, and 20% FDR levels. At 1% FDR, all RESET methods have negligible power, although this is consistent with many of the other methods. Interestingly, many of the finite-FDR controlling procedures such as the various RESET and AdaPT methods generally have better power than the asymptotic approaches, ZAP and AdaFDR, suggesting perhaps a misspecification of the fitted model used by these approaches. Figure 10 also shows the estimated FDR of each method. We find that in (a), AdaFDR’s estimated FDR at 20% (21.2% estimated FDR with 50% power) slightly exceeds the FDR threshold. This is possibly explained by the fact that AdaFDR only guarantees asymptotic FDP control. Consisitently, Chao and Fithian [8] reported inflated estimated FDRs at high thresholds in simulation studies for AdaFDR.
@@ 6.1.2 Simulation 6: p-value based testing with 100-dimensional side information
In Simulation 2 of Lei and Fithian [33], the side information for $m=2000$ hypotheses consists of 100 i.i.d draws from $U[0,1]$. This setup simulates ‘noisy’ side information since only the first 2 of the 100 draws are used in distinguishing between the true and false null hypotheses. Each p-value $p_{i}$ is drawn i.i.d from a two-group beta mixture model, where the true nulls have a uniform distribution and the false nulls have an alternative beta distribution. Specifically, the beta mixture model is given by:
@@
where $\mathbf{x}_{i}\in[0,1]^{100}$ denotes the side information of the $i$th hypothesis, $\pi(\mathbf{x}_{i})$ denotes proportion of the false nulls in the mixture, $h$ denotes the false null distribution of $Beta(1/\mu(\mathbf{x}_{i}),1)$. The $\pi(\mathbf{x}_{i})$’s and $\mu(\mathbf{x}_{i})$’s are determined by the following relationships with the 100-dimesional side information:
@@
There were a couple of tools we had to omit from the analysis of this data. Suprisingly, we found that AdaFDR was too slow in this simuation. Similarly, we were not able to apply AdaPT GAM or AdaPTg GAM as we did in the previous simulation, since it would have required using a smoothing spline on all 100 side-information variables, which is impractical. Hence, we only applied the GLMnet variants of AdaPT.
@@
Notably, the performance of AdaPT in these last two simulation setups is highly dependent on the choice of implementation. In Simulation 5, AdaPT GAM significantly outperforms AdaPT GLMnet while in Simulation 6, AdaPT GLMnet is superior (and AdaPT GAM is not practical). In contrast, RESET Ensemble selects from a collection of flexible machine learning models to adapt itself to the problem at hand.

### v2

@@ 5.2 Application to peptide detection
In this section we compare RESET to Adaptive Knockoffs in the peptide detection context using mass spectrometry data and show that (a) the Adaptive Knockoff methods are often computationally infeasible, (b) feasibility aside, RESET is more powerful than Adaptive Knockoff’s default filter, AdaKO GAM, when controlling the FDR, and (c) when controlling the FDP, RESET is able to typically discover more peptides than the vanilla FDP-SD. We provide a heuristic justification of Assumption 3, which is required for RESET’s theoretical guarantees, in Section S.15.
@@ 5.2.1 HEK293 data
We show that the application of Adaptive Knockoffs is often too slow to be used effectively in scenarios such as the peptide detection problem that involve a large amount of data. We demonstrate this using HEK293 data [9], a popular dataset used in benchmarking the development of new software [30, 50, 32] that we downloaded from the Proteomics Identification Database (PRIDE) [41]. After data preparation and searching the combined 24 HEK293 spectrum files, we determined the scores $Z_{i}$ for each target peptide and $\tilde{Z}_{i}$ for each decoy peptide using the XCorr score function (see Supplementary Section S.12 for details). Next, we competed each pair by recording the score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ and label $L_{i}=1$ if $Z_{i}>\tilde{Z}_{i}$ and $L_{i}=-1$ if $Z_{i}<\tilde{Z}_{i}$ (ties were randomly broken). In addition to its primary score we associate with each peptide-spectrum match (PSM) a set of features, such as the charge state or peptide length. Each winning peptide inherits these features, or side information $\mathbf{x}_{i}$, from its maximally associated PSM, allowing us to define the desired triples $(L,W,\mathbf{x})$. A complete description of the side information used is given in Table S4 of the Supplement.
@@ 5.2.2 PRIDE data
Due to the aforementioned runtime problem, we had to use small datasets when evaluating the power of Adaptive Knockoffs in the peptide detection context. Specifically, we focused on thirteen spectrum files (each with $<$ 33K PSMs) from the twenty PRIDE-20 datasets that we recently used in [18]. In the following results, we only considered searching the spectrum files once against a single target-decoy database, since unlike RESET we can only feasibly apply a single application of Adaptive Knockoffs. Data preparation, search settings and the detailed definition of the triples $(L,W,\mathbf{x})$ are given in Supplementary Section S.14 along with the list of side information used in Supplementary Table S5.555The side information here slightly differs from above due to different search settings.
@@
Figure 5: Comparison of RESET Ensemble and AdaKO GAM. For each of the 13 PRIDE datasets, we computed the average number of discoveries using 10 applications of RESET Ensemble, varying RESET’s internal seed. The figure shows a boxplot of those RESET Ensemble averages divided by the number of discoveries using AdaKO GAM (left boxplot) and TDC (right boxplot) using the 13 PRIDE-20 datasets (there are total of 13 points for each boxplot).
@@
Figure 5 shows the average number of discoveries reported using 10 applications of RESET Ensemble, each applied at 1% FDR to the same dataset, divided by the number of discoveries reported by AdaKO GAM (left boxplot) and TDC (right boxplot). In each of the 10 applications of RESET, we varied the internal seed which randomly splits the decoys into their training and estimating sets. Clearly, RESET Ensemble is overall more powerful than AdaKO GAM and TDC. Interestingly, the relative ratios of the number of discoveries between AdaKO GAM and TDC suggest that AdaKO GAM is reporting fewer discoveries than the generic TDC.
@@
In a second comparison, we looked at the performance of RESET Ensemble versus AdaKO RF because in this setup it was the most powerful implementation of Adaptive Knockoffs. In Figure 6A, the left boxplot displays the average number of discoveries by RESET Ensemble, divided by the number of discoveries by AdaKO RF. In six of the thirteen PRIDE spectrum files, we found RESET Ensemble to be the superior method, while in the remaining seven spectrum files, AdaKO RF produced more discoveries. In one dataset, AdaKO RF discovered 1217 peptides while RESET Ensemble only discovered 1104 peptides, about a 10% increase over RESET Ensemble.
@@
We also assessed the computational times of AdaKO RF in Figure 6B, using both the actual and estimated runtimes as outlined in Section S.13. Reassuringly, our estimated times were less than the actual times taken by each method to complete, thus validating the conservative nature of our estimated times in the previous section. Although we find that, when analyzing these smaller spectrum sets, AdaKO RF is the most powerful method of Adaptive Knockoffs, and is competitive to RESET Ensemble power-wise, it is again clear that AdaKO RF is generally too slow with some runtimes exceeding multiple days even for these small datasets. In fact, on the dataset where AdaKO RF outperforms RESET Ensemble by about 10%, it takes 3.07 days for AdaKO RF to discover its peptides while RESET Ensemble takes about half a minute to complete. Figure 6C shows that in the worst case, RESET takes roughly 1 minute. The issue of runtimes is especially problematic considering the fact that multiple spectrum files are often analyzed jointly, as in our HEK293 analysis in the previous section.
@@ 5.2.3 FDP control
Lastly, we evaluated RESET Ensemble’s FDP control using the HEK293 and all twenty PRIDE-20 spectrum files from the previous sections. As in Section 5.2.1, the joint HEK293 spectrum files were searched against 5 distinct target-decoy databases, where each target-decoy database was generated by concatenating the target database to a decoy database consisting of shuffled target peptides. Similarly, each PRIDE-20 spectrum file was searched against 10 distinct target-decoy databases. We used the coinflip version of FDP-SD [40], which provides a uniform increase in the number of discoveries (at the cost of some randomness in the discovery list, see Supplementary Algorithm S2). Our results using RESET Ensemble and FDP-SD are given in Table 3 and Figure 7.
@@
Analyzing the HEK293 results first, RESET Ensemble is clearly more powerful, with an approximately 9–10% increase in the average number of discoveries over FDP-SD at all confidence levels. Analyzing the PRIDE-20 datasets, while RESET Ensemble appears to be generally more powerful, in a few datasets it fails to make any discoveries at high confidence levels. Upon a closer look, the numbers of discoveries made by FDP-SD in these datasets are relatively small, with the largest average number of discoveries among these datasets being 417. When the number of discoveries is expected to be small and the confidence parameter is high, RESET may yield less power, even if the side information is reasonably informative — a phenomenon we dwell on in the Discussion. Regardless, more often than not RESET Ensemble produces more discoveries, and if multiple spectrum files are analyzed jointly, as we do with the HEK293 spectrum files, then the risk of having too few discoveries is reduced.
Figure 7: Comparison of RESET Ensemble and FDP-SD. The left panel shows boxplots of the average number of discoveries using RESET divided by the average number of discoveries using FDP-SD for each of the twenty PRIDE-20 datasets at a 1% FDP threshold and varying confidence parameters. The averages are taken over 10 applications, one for each of the 10 target-decoy databases used for each PRIDE-20 dataset. The right panel is a ‘zoomed’ in version of the left panel.
The results above consider a canonical ‘narrow’ search. Specifically, during the searching phase, each database peptide and spectrum are only allowed to match if the mass of the database peptide and the observed mass of the sample peptide generating the spectrum, the ‘spectrum mass’, differ by a small amount. Recently, ‘open’ searching, where each spectrum and peptide can match even if their masses differ by hundreds of Daltons, has become a popular method for detecting peptides [30]. Such a search allows for the detection of peptides whose masses have been altered by post-translational modifications (PTMs). We provide further detail in Supplementary Section S.14. In this setting, side information variables such as dm, which records the mass difference between the peptide and the spectrum, can be highly informative given that some differences will correspond to masses of common PTMs. Thus, it is not surprising that RESET, which can take advantage of such side information, improves on FDP-SD even further across all confidence levels in this context (Figure 8).
@@ 6.1 Numerical experiments
In this section, we consider the p-value setting, comparing RESET with the FDR controlling procedures we reviewed in Section 3.1 on a set of simulations introduced by Lei and Fithian [33]. As in the competition setting, each setup is designed to satisfy Assumption 3, which is required for our theoretical guarantees. Details of each method used are given in Supplementary Sections S.6S.10.
@@ 6.1.1 Simulation 5: p-value based testing with geometric side information
In this section, we look at Simulation 1 of Lei and Fithian [33]. For each hypothesis, they obtain a p-value $p_{i}$ from a one-sided normal test so that $p_{i}=1-\Phi(z_{i})$ where $z_{i}\sim\mathcal{N}(\mu,1)$ with $\mu=2$ for a false null hypothesis and $\mu=0$ for a true null hypothesis. A total of $m=2500$ hypotheses are used. Similar to Simulation 3 of Section 5.1.3, each hypothesis’ side information is defined as a unique pair of values $\mathbf{x}_{i}=\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ from $50\times 50$ equally spaced values in the square $\left[-100,100\right]\times\left[-100,100\right]$. The false null hypotheses correspond to certain regions in the square. They considered three scenarios: (a) when the false nulls form a circle in the middle of the square, (b) when the false nulls form a circle in the top right of the square and (c) when the false nulls form an ellipse (Figure 9). We used 100 independent runs, where in each run we randomly draw the 2500 p-values as described.
@@
Figure 9: Location of false null hypotheses according to the pair of values $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the square $\left[-100,100\right]\times\left[-100,100\right]$. The same picture is given in [33].
@@
Figure 10: RESET vs. other methods using two-dimensional side information with p-values. The first row plots the power and the second row plots the estimated FDR for each FDR threshold ranging from 1% to 20%. Each color and shape combination corresponds to a unique method. The three columns correspond to the hypotheses’ relationship to the side information according to Figure 9. The horizontal segments in the second row of panels mark the corresponding FDR thresholds on the y-axis. For readability, each point is jittered in the horizontal direction.
@@
Figure 10 shows the results of RESET compared with several p-value based methods that we reviewed in Section 3.1. We find that in the first two simulations, (a) and (b), RESET Ensemble is roughly equal to the best method in terms of power at the 5%, 10% and 20% FDR levels, along with AdaPT/AdaPTg GAM. However, in (c), clearly AdaPT/AdaPTg GAM are more powerful than the RESET methods, with RESET Ensemble coming arguably as third in the power ranking at the 5%, 10%, and 20% FDR levels. At 1% FDR, all RESET methods have negligible power, although this is consistent with many of the other methods. Interestingly, many of the finite-FDR controlling procedures such as the various RESET and AdaPT methods generally have better power than the asymptotic approaches, ZAP and AdaFDR, suggesting perhaps a misspecification of the fitted model used by these approaches. Figure 10 also shows the estimated FDR of each method. We find that in (a), AdaFDR’s estimated FDR at 20% (21.2% estimated FDR with 50% power) slightly exceeds the FDR threshold. This is possibly explained by the fact that AdaFDR only guarantees asymptotic FDP control. Consistently, Chao and Fithian [8] reported inflated estimated FDRs at high thresholds in simulation studies for AdaFDR.
@@ 6.1.2 Simulation 6: p-value based testing with 100-dimensional side information
In Simulation 2 of Lei and Fithian [33], the side information for $m=2000$ hypotheses consists of 100 i.i.d. draws from $U[0,1]$. This setup simulates ‘noisy’ side information since only the first 2 of the 100 draws are used in distinguishing between the true and false null hypotheses. Each p-value $p_{i}$ is drawn i.i.d. from a two-group beta mixture model, where the true nulls have a uniform distribution and the false nulls have an alternative beta distribution. Specifically, the beta mixture model is given by:
@@
where $\mathbf{x}_{i}\in[0,1]^{100}$ denotes the side information of the $i$th hypothesis, $\pi(\mathbf{x}_{i})$ denotes proportion of the false nulls in the mixture, $h$ denotes the false null distribution of $Beta(1/\mu(\mathbf{x}_{i}),1)$. The $\pi(\mathbf{x}_{i})$’s and $\mu(\mathbf{x}_{i})$’s are determined by the following relationships with the 100-dimensional side information:
@@
There were a couple of tools we had to omit from the analysis of this data. Surprisingly, we found that AdaFDR was too slow in this simulation. Similarly, we were not able to apply AdaPT GAM or AdaPTg GAM as we did in the previous simulation, since it would have required using a smoothing spline on all 100 side-information variables, which is impractical. Hence, we only applied the GLMnet variants of AdaPT.
@@
Notably, the performance of AdaPT in these last two simulation setups is highly dependent on the choice of implementation. In Simulation 5, AdaPT GAM significantly outperforms AdaPT GLMnet while in Simulation 6, AdaPT GLMnet is superior and AdaPT GAM is not even practical. In contrast, RESET Ensemble selects from a collection of flexible machine learning models to adapt itself to the problem at hand.